Starling Co. is considering disposing of a machine with a book value of $21,100 and estimated remaining life of five years. The old machine can be sold for $5,200. A new high-speed machine can be purchased at a cost of 65,200. It will have a useful life of five years and no residual value. It is estimated that the annual variable manufacturing costs will be reduced from $23,500 to $19,100 if the new machine is purchased. The differential effect on income for the new machine for the entire five years is a. increase of $49,400 b. decrease of $49,400 c. increase of $38,000 d. decrease of $38,000
Answer #1

Calculate differential effect

Retain Replace
New machine 65200
Variable cost 23500*5 = 117500 19100*5 = 95500
Scrap value -5200
Total relevant cost 117500 155500

Net income decrease by = 155500-117500 = 38000

So answer is d) Decrease of $38000